OpenStudy (anonymous):

Simplify the quotient 4.5×10^3/9×10^7 write your answer in scientific notation SHOW EVERY DETAIL

OpenStudy (nincompoop):

what is \(\frac{4.5}{9} \) in simplest form?

OpenStudy (nincompoop):

laws of exponent \(\huge \frac{p^n}{p^m} = p^{n-m} \)

OpenStudy (nincompoop):

thank you for your time

OpenStudy (anonymous):

Sorry my internet went out

OpenStudy (anonymous):

And s 1 I think

OpenStudy (anonymous):

@nnesha

OpenStudy (anonymous):

10^10?